# and Bethe Center for Theoretical Physics, Universität Bonn, 53115 Bonn, Germany

### 20 Feb 2024 v1

@@ 1 Introduction
In recent decades, the Lüscher method Luscher:1990ux has become a standard tool for the extraction of the scattering phase shifts from the finite-volume energy levels, measured in lattice QCD. The method has been generalized to the case of moving frames, particles with spin and coupled two-body channels (see here Rummukainen:1995vs ; Lage:2009zv ; Bernard:2010fp ; He:2005ey ; Liu:2005kr ; Hansen:2012tf ; Briceno:2012yi ; Li:2012bi ; Guo:2012hv ; Leskovec:2012gb ; Kim:2005gf ; Gockeler:2012yj for a representative list of references). In all cases, the formalism is based on fundamental assumptions, namely:
  • The interactions between particles are short-range. The relation $R/L\ll 1$ holds, where $R$ is the characteristic range of interaction and $L$ denotes a size of the cubic box (the spatial extension of the lattice) in which the system is placed. The quantity $R$ is typically given by the inverse of the lightest mass in the theory, $R\sim M^{-1}$.
@@
For completeness, we also note that, to the best of our knowledge, Ref. Beane:2014qha is the only place in the literature where the two-body quantization condition in the presence of the Coulomb force has been considered. The latter is, however, treated perturbatively, and only the non-derivative strong coupling is taken into account. In the following we shall see that including the derivative couplings is a rather subtle issue.
@@ 2.2 NREFT framework
The phase shift is related to the on-shell partial wave amplitudes via:
@@ 3.2.1 Negative energies, $q_{0}^{2}<0$
$+$ $4\pi\sum_{\ell^{\prime\prime}m^{\prime}}\int_{0}^{\infty}\frac{k^{2}dk}{(2\pi)^{3}}\,V_{L}^{\ell}(p,k)\frac{f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(k)-f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(-k)}{k^{2}-q_{0}^{2}}\,T_{L}^{\ell^{\prime\prime}m^{\prime\prime},\ell^{\prime}m^{\prime}}(p,q;q_{0}^{2})\,,$
@@
The quantity $f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(k)$ is analytic in the upper half-plane of the variable $k$.
@@
$T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(p,q;q_{0}^{2})=T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(-p,q;q_{0}^{2})=T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(p,-q;q_{0}^{2})=T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(-p,-q;q_{0}^{2})\,.$
@@ 3.2.2 Positive energies, $q_{0}^{2}>0$
Neglecting exponentially suppressed contributions from the long-range interactions, one could reduce the calculation of the function $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$ to the solution of the system of linear equations in the angular momentum basis. This equation has the following form:
@@ 3.4 Exponentially suppressed effects
Up to now, we have consistently dropped the exponentially suppressed effects. However, as mentioned already in the introduction, these effects can turn relatively large, owing to the small mass scale. In the case of, say, $NN$ scattering, one may indirectly estimate the size of the exponential effects, comparing the finite-volume spectra obtained in the plane-wave basis with the solutions of the modified Lüscher equation with the same input. A simpler method to estimate the size of the exponential effects is the comparison of the modified Lüscher functions $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$, calculated in the plane wave basis and in the angular momentum basis. This comparison does not involve any parameters that characterize the short-range interactions.
@@ 4 Comparison with the existing approaches
In the recent paper Raposo:2023oru a modified two-body quantization condition has been derived in the presence of both the long- and short-range forces. The authors present their central result in two different forms. Namely, Eq. (3.63) of that paper is written down in a plane wave and angular momentum basis. From this point of view, it bears strong resemblance with the approach of Ref. Meng:2021uhz , however, with a conceptual difference. Namely, all short-range interactions in Ref. Raposo:2023oru are summed up and enter the quantization condition through an auxiliary on-shell $K$-matrix $\bar{K}^{\sf os}$. No particular parameterization of the on-shell $K$-matrix $\bar{K}^{\sf os}$ is specified. We note, however, that for any realistic application to lattice data involving partial wave mixing such a parameterization would be required. In contrast, the long- and short-range interactions are treated on equal footing in Ref. Meng:2021uhz , and the short-range $K$-matrix is implicitly parameterized in terms of the effective couplings appearing in the Hamiltonian.
Furthermore, in Eqs. (5.3) and (5.4) of Ref. Raposo:2023oru the authors recast their central result in the angular momentum basis. This result bears close analogy to our modified quantization condition. For example, the quantity $F^{T}$ from their Eq. (5.4) is similar to our modified Lüscher function $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$. The main difference, as already noted in the introduction, is that the authors of Ref. Raposo:2023oru propose a two-step procedure for the analysis of lattice data. Namely, at the first step, an auxiliary matrix $\bar{K}^{\sf os}$ is determined from data. At the next step, $\bar{K}^{\sf os}$ is substituted into the integral equation which is solved to obtain the physical $K$-matrix. We propose to unite these two steps in one – in our approach, the auxiliary $K$-matrix is related to the physical one at the same CM energy through a simple algebraic expression.
@@
    Technical convenience. The quantization condition can be written down in the plane wave basis as well as in the angular momentum basis. Given modern computing capacities, the difference between these two representations is not a decisive factor anymore. Despite this, we still prefer a more compact representation in the angular momentum basis, which reduces to a single algebraic equation if the partial-wave mixing for the short-range interactions can be neglected. The same statement applies to relating physical observables to quantities extracted from the fit to lattice data. For example, in order to relate $\bar{K}^{\sf os}$ to the physical $K$-matrix, integral equations need to be solved Raposo:2023oru , whereas the corresponding link in our approach is given by a simple algebraic relation.
@@ 5 Conclusions
    Note that in this paper we deliberately ignored all issues related with the spin of particles, moving frames, relativistic effects, etc. All this in inessential in the context of the problems considered here and would only blur the discussion.
@@
    It remains to be seen, whether the Coulomb interaction can be treated consistently in the same manner, and whether the results would add something substantial to the findings of Ref. Beane:2014qha . Here, it should be also mentioned that, due to the removal of the zero mode of the Coulomb field in case of periodic boundary conditions, the resulting Lagrangian is not local anymore. This, in its turn, might cause problems in the matching of the non-relativistic effective field theory, which is used for the derivation of the Lüscher equation, to its relativistic counterpart (see, e.g., Ref. Davoudi:2018qpl ). In this context, it would be interesting to explore the possibility of using different boundary conditions.
@@
Acknowledgments: The authors thank Vadim Baru, Evgeny Epelbaum, Jambul Gegelia, Christoph Hanhart, Bai Long, Lu Meng and Fernando Romero-Lopez for interesting discussions. The work of R.B, F.M. and A.R. was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 196253076 – TRR 110 and by the Ministry of Culture and Science of North Rhine-Westphalia through the NRW-FAIR project. A.R., in addition, thanks Volkswagenstiftung (grant no. 93562) and the Chinese Academy of Sciences (CAS) President’s International Fellowship Initiative (PIFI) (grant no. 2024VMB0001) for the partial financial support. The work of J.-Y.P. and J.-J.W. was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12135011, 12175239, 12221005, and by the National Key R&D Program of China under Contract No. 2020YFA0406400. H.-W.H. was supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundati on) – Project ID 279384907 – SFB 1245.

### 28 Feb 2024 v2

@@ 1 Introduction
In recent decades, the Lüscher method Luscher:1990ux has become a standard tool for the extraction of the scattering phase shifts from the finite-volume energy levels, measured in lattice QCD. The method has been generalized to the case of moving frames, particles with spin and coupled two-body channels (see here Rummukainen:1995vs ; Lage:2009zv ; Bernard:2010fp ; He:2005ey ; Liu:2005kr ; Hansen:2012tf ; Briceno:2012yi ; Li:2012bi ; Guo:2012hv ; Leskovec:2012gb ; Kim:2005gf ; Gockeler:2012yj for a representative list of references). In all cases, the formalism is based on the fundamental assumptions, namely:
  • The interactions between particles are short-range. The relation $R/L\ll 1$ holds, where $R$ is the characteristic range of interaction and $L$ denotes the size of the cubic box (the spatial extension of the lattice) in which the system is placed. The quantity $R$ is typically given by the inverse of the lightest mass in the theory, $R\sim M^{-1}$.
@@
For completeness, we also note that, the two-body quantization condition (as well as the two-body Lellouch-Lüscher formula) in the presence of the Coulomb force was considered in Refs. Beane:2014qha ; Cai:2018why ; Christ:2021guf . In the final expressions, the Coulomb potential has been treated perturbatively in the fine structure constant. Such a treatment can be justified for sufficiently small values of the box size $L$.
@@ 2.2 NREFT framework
The phase shift is related to the on-shell partial-wave amplitudes via:
@@ 3.2.1 Negative energies, $q_{0}^{2}<0$
$+$ $4\pi\sum_{\ell^{\prime\prime}m^{\prime}}\int_{0}^{\infty}\frac{k^{2}dk}{(2\pi)^{3}}\,V_{L}^{\ell}(p,k)\frac{f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(k)+f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(-k)}{k^{2}-q_{0}^{2}}\,T_{L}^{\ell^{\prime\prime}m^{\prime\prime},\ell^{\prime}m^{\prime}}(p,q;q_{0}^{2})\,,$
@@
The quantity $f_{\ell m,\ell^{\prime\prime}m^{\prime\prime}}(k)$ is analytic in the upper half-plane of the variable $k$ and vanishes exponentially, when ${\rm Im}\,k\to+\infty$. Note that $\ell+\ell^{\prime}$ is always even for identical particles.
@@
$T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(p,q;q_{0}^{2})$ $=$ $(-1)^{\ell}T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(-p,q;q_{0}^{2})=(-1)^{\ell^{\prime}}T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(p,-q;q_{0}^{2})$ (73)
$=$ $T_{L}^{\ell m,\ell^{\prime}m^{\prime}}(-p,-q;q_{0}^{2})\,.$
@@ 3.2.2 Positive energies, $q_{0}^{2}>0$
Neglecting exponentially suppressed contributions from the long-range interactions, one could reduce the calculation of the function $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$ to the solution of the system of linear equations in the angular-momentum basis. This equation has the following form:
@@ 3.4 Exponentially suppressed effects
Up to now, we have consistently dropped the exponentially suppressed effects. However, as mentioned already in the introduction, these effects can turn relatively large, owing to the small mass scale. In the case of, say, $NN$ scattering, one may indirectly estimate the size of the exponential effects, comparing the finite-volume spectra obtained in the plane-wave basis with the solutions of the modified Lüscher equation with the same input. A simpler method to estimate the size of the exponential effects is the comparison of the modified Lüscher functions $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$, calculated in the plane-wave basis and in the angular-momentum basis. This comparison does not involve any parameters that characterize the short-range interactions.
@@ 4 Comparison with the existing approaches
In the recent paper Raposo:2023oru a modified two-body quantization condition has been derived in the presence of both the long- and short-range forces. The authors present their central result in two different forms. Namely, Eq. (3.63) of that paper is written down in a plane wave and angular-momentum basis. From this point of view, it bears strong resemblance with the approach of Ref. Meng:2021uhz , however, with a conceptual difference. Namely, all short-range interactions in Ref. Raposo:2023oru are summed up and enter the quantization condition through an auxiliary on-shell $K$-matrix $\bar{K}^{\sf os}$. No particular parameterization of the on-shell $K$-matrix $\bar{K}^{\sf os}$ is specified. We note, however, that for any realistic application to lattice data involving partial-wave mixing such a parameterization would be required. In contrast, the long- and short-range interactions are treated on equal footing in Ref. Meng:2021uhz , and the short-range $K$-matrix is implicitly parameterized in terms of the effective couplings appearing in the Hamiltonian.
Furthermore, in Eqs. (5.3) and (5.4) of Ref. Raposo:2023oru the authors recast their central result in the angular-momentum basis. This result bears close analogy to our modified quantization condition. For example, the quantity $F^{T}$ from their Eq. (5.4) is similar to our modified Lüscher function $H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})$. The main difference, as already noted in the introduction, is that the authors of Ref. Raposo:2023oru propose a two-step procedure for the analysis of lattice data. Namely, at the first step, an auxiliary matrix $\bar{K}^{\sf os}$ is determined from data. At the next step, $\bar{K}^{\sf os}$ is substituted into the integral equation which is solved to obtain the physical $K$-matrix. We propose to unite these two steps in one – in our approach, the auxiliary $K$-matrix is related to the physical one at the same CM energy through a simple algebraic expression.
@@
    Technical convenience. The quantization condition can be written down in the plane-wave basis as well as in the angular-momentum basis. Given modern computing capacities, the difference between these two representations is not a decisive factor anymore. Despite this, we still prefer a more compact representation in the angular-momentum basis, which reduces to a single algebraic equation if the partial-wave mixing for the short-range interactions can be neglected. The same statement applies to relating physical observables to quantities extracted from the fit to lattice data. For example, in order to relate $\bar{K}^{\sf os}$ to the physical $K$-matrix, integral equations need to be solved Raposo:2023oru , whereas the corresponding link in our approach is given by a simple algebraic relation.
@@ 5 Conclusions
    Note that in this paper we deliberately ignored all issues related with the spin of particles, moving frames, relativistic effects, etc. All this is inessential in the context of the problems considered here and would only blur the discussion.
@@
    It remains to be seen, whether the Coulomb interaction can be treated consistently in the same manner, and whether the results would add something substantial to the findings of Refs. Beane:2014qha ; Cai:2018why ; Christ:2021guf . Here, it should be also mentioned that, due to the removal of the zero mode of the Coulomb field in case of periodic boundary conditions, the resulting Lagrangian is not local anymore. This, in its turn, might cause problems in the matching of the non-relativistic effective field theory, which is used for the derivation of the Lüscher equation, to its relativistic counterpart (see, e.g., Ref. Davoudi:2018qpl ). In this context, it would be interesting to explore the possibility of using different boundary conditions. An alternative to this would be to use the formulation with massive protons, see, e.g. Endres:2015gda .
@@
Acknowledgments: The authors thank Vadim Baru, Evgeny Epelbaum, Jambul Gegelia, Christoph Hanhart, Nils Hermanssohn-Truedsson, Bai-Long Hoid, Lu Meng and Fernando Romero-Lopez for interesting discussions. The work of R.B, F.M. and A.R. was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 196253076 – TRR 110 and by the Ministry of Culture and Science of North Rhine-Westphalia through the NRW-FAIR project. A.R., in addition, thanks Volkswagenstiftung (grant no. 93562) and the Chinese Academy of Sciences (CAS) President’s International Fellowship Initiative (PIFI) (grant no. 2024VMB0001) for the partial financial support. The work of J.-Y.P. and J.-J.W. was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12135011, 12175239, 12221005, and by the National Key R&D Program of China under Contract No. 2020YFA0406400, and by the Chinese Academy of Sciences under Grant No. YSBR-101, and by the Xiaomi Foundation / Xiaomi Young Talents Program. H.-W.H. was supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundati on) – Project ID 279384907 – SFB 1245.